OpenStudy (anonymous):

A 10000 N vehicle is stalled 1/4 of the way across a bridge. The bridge is 20m in length, from support to support. The bridges weight of 50,000N is uniformly distributed. <-------20m-----> Fa---(Car)--------Fb <--5m--> a) find the support force of Fb b) find the support force of Fa

OpenStudy (fretje):

a: supposing this is no space bridge or something on high altitude we can analyse the forces equation by taking a point on the bridge (I choose point a (where Fa is placed)). and I calculate the momenta around that point of the forces. The sum of the moments should be zero as the construction is in statical balance, and point a does not take up any moments, only translational forces. The car makes a moment clockwise (clockwise is +) of 10000N*5m = 50000Nm The point b of the bridge should therefore make the same moment of -50000Nm The force Fcb is therefore -50000Nm/20m = 2500 Newton Than I add half off the bridges weight : 25000N Fb = 27500N b: to get the now I take the the point b for my balance equation in point b we have following moments: Mb = 0 = Fca*20m - 10000N*15m Solve for Fca = 150000N/20 = 7500N I add half of the bridges weight and get Fa = 57500N

OpenStudy (fretje):

ok a bit fuzzy, but the priciple is that one can calculate first the forces on the points a and b from one weight (ex the car), just like there are no other forces, then the forces of the other weight (the bridge), and add them.(superposition of forces)